# A Survey Analyzing Generalization in Deep Reinforcement Learning

### 04 Jan 2024 v1

@@ Abstract
Reinforcement learning research obtained significant success and attention with the utilization of deep neural networks to solve problems in high dimensional state or action spaces. While deep reinforcement learning policies are currently being deployed in many different fields from medical applications to self driving vehicles, there are still ongoing questions the field is trying to answer on the generalization capabilities of deep reinforcement learning policies. In this paper, we will outline the fundamental reasons why deep reinforcement learning policies encounter overfitting problems that limit their robustness and generalization capabilities. Furthermore, we will formalize and unify the diverse solution approaches to increase generalization, and overcome overfitting in state-action value functions. We believe our study can provide a compact systematic unified analysis for the current advancements in deep reinforcement learning, and help to construct robust deep neural policies with improved generalization abilities.
@@ 1 Introduction
The performance of reinforcement learning algorithms has been boosted with the utilization of deep neural networks as function approximators (Mnih et al., 2015). Currently, it is possible to learn deep reinforcement learning policies that can operate in large state and/or action space MDPs Silver et al. (2017); Vinyals et al. (2019). This progress consequently resulted in building reasonable deep reinforcement learning policies that can play computer games with high dimensional state representations (e.g. Atari, StarCraft), solve complex robotics control tasks, design algorithms (Mankowitz et al., 2023; Fawzi et al., 2022), and play some of the most complicated board games (e.g. Chess, Go) (Schrittwieser et al., 2020). However, deep reinforcement learning algorithms also experience several problems caused by their overall limited generalization capabilities. Some studies demonstrated these problems via adversarial perturbations introduced to the state observations of the policy (Huang et al., 2017; Kos & Song, 2017; Korkmaz, 2022), several focused on exploring the fundamental issues with function approximation, estimation biases in the state-action value function (Hasselt et al., 2016), or with new architectural design ideas (Wang et al., 2016).
The fact that we are not able to completely explore the entire MDP for high dimensional state representation MDPs, even with deep neural networks as function approximators, is one of the root problems that limits generalization. On top of this, some portion of the problems are directly caused by the utilization of deep neural networks and thereby the intrinsic problems inherited from their utilization (Goodfellow et al., 2015; Szegedy et al., 2014; Korkmaz, 2022).
In this paper we will focus on generalization in deep reinforcement learning and the underlying causes of the limitations deep reinforcement learning research currently faces. In particular, we will try to answer the following questions:
@@
  • What has been done to overcome the overfitting problems that deep reinforcement learning algorithms have encountered so far?
  • What future directions are there for reinforcement learning research to obtain higher level generalization abilities for deep neural policies?
To answer these questions we will go through research connecting several subfields in reinforcement learning on the problems and corresponding proposed solutions regarding generalization. In this paper we introduce a categorization of the different methods used to both achieve and test generalization, and use it to systematically summarize and consolidate the current body of research. We further describe the issue of value function overestimation, and the role of exploration in overfitting in reinforcement learning. Furthermore, we explain new emerging research areas that can potentially target these questions in the long run including meta-reinforcement learning and lifelong learning. We hope that our paper can provide a compact overview and unification of the current advancements and limitations in the field.
@@ 2 Preliminaries on Deep Reinforcement Learning
The aim in deep reinforcement learning is to learn a policy via interacting with an environment in a Markov Decision Process (MDP) that maximize expected cumulative discounted rewards. An MDP is represented by a tuple $\mathcal{M}=(S,A,P,r,\rho_{0},\gamma)$, where $S$ represents the state space, $A$ represents the action space, $r:S\times A\to\mathbb{R}$ is a reward function, $\mathcal{P}:S\times A\to\Delta(S)$ is a transition probability kernel, $\rho_{0}$ represents the initial state distribution, and $\gamma$ represents the discount factor. The objective in reinforcement learning is to learn a policy $\pi:S\to\Delta(A)$ which maps states to probability distributions on actions in order to maximize the expected cumulative reward $R=\mathbb{E}\sum_{t=0}^{T-1}\gamma^{t}r(s_{t},a_{t})$ where $a_{t}\sim\pi(s_{t}),s_{t+1}\sim\mathcal{P}(s_{t},a_{t})$. In $Q$-learning the goal is to learn the optimal state-action value function (Watkins, 1989)
$Q^{*}(s,a)=R(s,a)+\sum_{s^{\prime}\in S}P(s^{\prime}|s,a)\max_{a^{\prime}\in A}Q^{*}(s^{\prime},a^{\prime}).$ (1)
This is achieved via iterative Bellman update which updates $Q(s_{t},a_{t})$ by
@@
Thus, the optimal policy is determined by choosing the action $a^{*}(s)=\operatorname*{arg\,max}_{a}Q(s,a)$ in state $s$. In high dimensional state space or action space MDPs the optimal policy is decided via a function-approximated state-action value function represented by a deep neural network. In a parallel line of algorithm families the policy itself is directly parametrized by $\pi_{\theta}$, and the gradient estimator used in learning is
@@
where $Q(s_{t},a_{t})$ refers to the state-action value function at timestep $t$.
@@ 3 How to Achieve Generalization?
To be able to categorize different paths to achieve generalization first we will provide a definition meant to capture the behavior of a generic reinforcement learning algorithm.
Definition 3.1.
@@
A reinforcement learning training algorithm $\mathcal{A}$ learns a policy $\pi$ by interacting with an MDP $\mathcal{M}$. We divide up the execution of $\mathcal{A}$ into discrete time steps as follows. At each time $t$, the algorithm chooses a state $s_{t}$, takes an action $a_{t}$, observes a transition to state $s^{\prime}_{t}$ with corresponding reward $r_{t}=r(s_{t},a_{t},s^{\prime}_{t})$. We define the history of algorithm $\mathcal{A}$ in MDP $\mathcal{M}$ to be the sequence $H_{t}=(s_{0},a_{0},s^{\prime}_{0},r_{0}),\dots(s_{t},a_{t},s^{\prime}_{t},r_{t})$ of all the transitions observed by the algorithm so far. We require that state and action $(s_{t},a_{t})$ chosen at time $t$ are a function only of $H_{t-1}$, i.e the transitions observed so far by $\mathcal{A}$. At time $t=T$, the algorithm stops and outputs a policy $\pi$.
@@
Intuitively, a reinforcement learning algorithm performs a sequence of queries $(s_{t},a_{t})$ to the MDP, and observes the resulting state transitions and rewards. In order to be as generic as possible, the definition makes no assumptions about how the algorithm chooses the sequence of queries. Notably, if taking action $a_{t}$ in state $s_{t}$ leads to a transition to state $s^{\prime}_{t}$, there is no requirement that $s_{t+1}=s^{\prime}_{t}$. Indeed, the only assumption is that $(s_{t+1},a_{t+1})$ depends only on $H_{t}$, the history of transitions observed so far. This allows the definition to capture deep reinforcement learning algorithms, which may choose to query states and actions in a complex way based on previously observed state transitions. Based on this definition of generic reinforcement learning algorithm, we will now further define the different techniques proposed to achieve generalization.
@@
Let $\mathcal{A}$ be a training algorithm that takes as input an MDP and outputs a policy. Given an MDP $\mathcal{M}=(S,A,P,r,\rho_{0},\gamma)$, a rewards transforming generalization method $\mathcal{G}_{R}$ is given by a sequence of functions $F_{t}:(S\times A\times S\times\mathbb{R})^{t}\to\mathbb{R}$. The method attempts to achieve generalization by running $\mathcal{A}$ on MDP $\mathcal{M}$, but modifying the rewards at each time $t$ to be $\tilde{r}_{t}(s_{t},a_{t},s^{\prime}_{t})=F_{t-1}(H_{t-1})$, where $H_{t-1}$ is the history of algorithm $\mathcal{A}$ when running with the perturbed rewards.
In summary, a rewards transforming generalization methods simply runs the original algorithm, but modifies the observed rewards. Similarly, we define two additional generalization methods which run the original algorithm while modifying states and transition probabilities respectively.
@@
Let $\mathcal{A}$ be a training algorithm that takes as input an MDP and outputs a policy. Given an MDP $\mathcal{M}=(S,A,P,r,\rho_{0},\gamma)$, a state transforming generalization method $\mathcal{G}_{S}$ is given by a sequence of functions $F_{t}:(S\times A\times S\times\mathbb{R})^{t}\times S\to S$. The method attempts to achieve generalization by running $\mathcal{A}$ on MDP $\mathcal{M}$, but modifying the state chosen at time $t$ to be $\tilde{s}_{t}=F_{t-1}(H_{t-1},s_{t})$, where $H_{t-1}$ is the history of algorithm $\mathcal{A}$ when running with the perturbed states.
@@
Let $\mathcal{A}$ be a training algorithm that takes as input an MDP and outputs a policy. Given an MDP $\mathcal{M}=(S,A,P,r,\rho_{0},\gamma)$, a transition probability transforming generalization method $\mathcal{G}_{\mathcal{P}}$ is given by a sequence of functions $F_{t}:(S\times A\times S\times\mathbb{R})^{t}\times(S\times A\times S)\to\mathbb{R}$. The method attempts to achieve generalization by running $\mathcal{A}$ on MDP $\mathcal{M}$, but modifying the transition probabilities at time $t$ to be $\tilde{P}(s_{t},a_{t},s^{\prime}_{t})=F_{t-1}(H_{t-1},s_{t},a_{t},s^{\prime}_{t})$, where $H_{t-1}$ is the history of algorithm $\mathcal{A}$ when running with the perturbed transition probabilities.
The last type of generalization method we define is based on directly modifying the way in which the training algorithm chooses the state and action pair for the next time step. While this definition is broad enough to capture very complex changes to the training algorithm, in practice the choice of modification generally has a simple description.

### 30 Oct 2024 v2

@@ Abstract
Reinforcement learning research obtained significant success and attention with the utilization of deep neural networks to solve problems in high dimensional state or action spaces. While deep reinforcement learning policies are currently being deployed in many different fields from medical applications to large language models, there are still ongoing questions the field is trying to answer on the generalization capabilities of deep reinforcement learning policies. In this paper, we will formalize and analyze generalization in deep reinforcement learning. We will explain the fundamental reasons why deep reinforcement learning policies encounter overfitting problems that limit their generalization capabilities. Furthermore, we will categorize and explain the manifold solution approaches to increase generalization, and overcome overfitting in deep reinforcement learning policies. From exploration to adversarial analysis and from regularization to robustness our paper provides an analysis on a wide range of subfields within deep reinforcement learning with a broad scope and in-depth view. We believe our study can provide a compact guideline for the current advancements in deep reinforcement learning, and help to construct robust deep neural policies with higher generalization skills.
@@ 1 Introduction
The performance of reinforcement learning algorithms (Watkins,, 1989; Sutton,, 1984, 1988) has been boosted with the utilization of deep neural networks as function approximators (Mnih et al.,, 2015). Currently, it is possible to learn deep reinforcement learning policies that can operate in large state and/or action space MDPs (Silver et al.,, 2017; Vinyals et al.,, 2019). This progress consequently resulted in building reasonable deep reinforcement learning policies that can play computer games with high dimensional state representations (e.g. Atari, StarCraft), solve complex robotics control tasks, design algorithms (Mankowitz et al.,, 2023; Fawzi et al.,, 2022), guide large language models (OpenAI,, 2023; Google Gemini,, 2023), and play some of the most complicated board games (e.g. Chess, Go) (Schrittwieser et al.,, 2020). However, deep reinforcement learning algorithms also experience several problems caused by their overall limited generalization capabilities. Some studies demonstrated these problems via adversarial perturbations introduced to the state observations of the policy (Huang et al.,, 2017; Kos and Song,, 2017; Korkmaz,, 2022; Korkmaz and Brown-Cohen,, 2023), several focused on exploring the fundamental issues with function approximation, estimation biases in the state-action value function (Thrun and Schwartz,, 1993; van Hasselt,, 2010), or with new architectural design ideas (Wang et al.,, 2016). The fact that we are not able to completely explore the entire MDP for high dimensional state representation MDPs, even with deep neural networks as function approximators, is one of the root problems that limits generalization. On top of this, some portion of the problems are directly caused by the utilization of deep neural networks and thereby the intrinsic problems inherited from their utilization (Goodfellow et al.,, 2015; Szegedy et al.,, 2014; Korkmaz,, 2022; Korkmaz, 2024b, ).
In order to address open questions on generalization in deep reinforcement learning, there needs to be some commonly agreed standard of what is meant by generalization. Currently, different aspects of generalization are considered in various subfields either working on the fundamental questions regarding or the applications of deep reinforcement learning. We take the point of view in this paper that these various aspects can, and should, be described and studied in a unified way. In particular, we argue that the various approaches to generalization can be succinctly classified based on which part of the Markov Decision Process is expected to vary. We make this classification formal and unify how much current work on generalization in deep reinforcement learning fits clearly into the classification we introduce. In this paper we will focus on generalization in deep reinforcement learning and the underlying causes of the limitations deep reinforcement learning research currently faces. In particular, we will try to answer the following questions:
  • How can we formalize the concept of generalization in deep reinforcement learning?
@@
  • What has been done to overcome the overfitting problems that deep reinforcement learning algorithms have encountered so far, and to enable deep neural policies to generalize to non-stationary complex environments?
To answer these questions we will go through research connecting several subfields in reinforcement learning on the problems and corresponding proposed solutions regarding generalization. In this paper we introduce a formal definition of generalization and categorization of the different methods used to both achieve and assess generalization, and use it to systematically summarize and consolidate the current body of research. We further describe the issue of value function overestimation, and the role of exploration in overfitting in reinforcement learning. Furthermore, we explain new emerging research areas that can potentially target these questions in the long run including meta-reinforcement learning and lifelong learning. The objective of the paper is to introduce a formal generalization definition and provide a compact overview and unification of the current advancements and limitations in the field.
@@ 2 Preliminaries on Deep Reinforcement Learning
The aim in deep reinforcement learning is to learn a policy via interacting with an environment in a Markov Decision Process (MDP) that maximize expected cumulative discounted rewards. An MDP is represented by a tuple $\mathcal{M}=(S,A,\mathcal{P},r,\rho_{0},\gamma)$, where $S$ represents the state space, $A$ represents the action space, $r:S\times A\to\mathbb{R}$ is a reward function, $\mathcal{P}:S\times A\to\Delta(S)$ is a transition probability kernel, $\rho_{0}$ represents the initial state distribution, and $\gamma$ represents the discount factor. The objective in reinforcement learning is to learn a policy $\pi:S\times A\to\mathbb{R}$ which maps states to probability distributions on actions in order to maximize the expected cumulative reward $R=\mathbb{E}\sum_{t=0}^{T-1}\gamma^{t}r(s_{t},a_{t})$ where $a_{t}\sim\pi(s_{t},\cdot),s_{t+1}\sim\mathcal{P}(\cdot|s_{t},a_{t})$. The temporal difference updates achieves this objective by updating the value function $V(s)$ (Sutton,, 1984, 1988)
$V(s_{t})\leftarrow V(s_{t})+\alpha[r(s_{t+1},a)+\gamma V(s_{t+1})-V(s_{t})]$ (1)
While Equation 1 represents the one-step temporal difference update, i.e. TD(0), it is further possible to consider multi-step TD which focuses on multi-step return, i.e. TD($\lambda$). In $Q$-learning the goal is to learn the optimal state-action value function (Watkins,, 1989)
$Q^{*}(s,a)=r(s,a)+\sum_{s^{\prime}\in S}\mathcal{P}(s^{\prime}|s,a)\max_{a^{\prime}\in A}Q^{*}(s^{\prime},a^{\prime}).$ (2)
This is achieved via iterative Bellman update (Bellman,, 1957; Bellman and Dreyfus,, 1959) which updates $Q(s_{t},a_{t})$ by
@@
Thus, the optimal policy is determined by choosing the action $a^{*}(s)=\operatorname*{arg\,max}_{a}Q(s,a)$ in state $s$. The optimal Bellman operator is (Bellman,, 1957)
$\mathcal{B}Q(s,a)\coloneqq\mathbb{E}[r(s,a)]+\gamma\mathbb{E}_{\mathcal{P}}[\max_{a^{\prime}}Q(s^{\prime},a^{\prime})]$
In high dimensional state space or action space MDPs the optimal policy is decided via a function-approximated state-action value function represented by a deep neural network. The loss function in deep reinforcement learning is the quadratic difference between that target network and the current state-action value function.
$\mathcal{L}_{i}(\theta_{i})=\mathbb{E}_{e\sim\mathcal{D}}[(r(s,a)+\gamma\max_{a^{\prime}}Q(s^{\prime},a^{\prime},\theta_{\textrm{target}})-Q(s,a,\theta_{i}))^{2}]$
where $\mathcal{D}$ is the experience replay buffer (Lin,, 1993) in which the experiences $e=\{s_{t},a_{t},r_{t},s_{t+1}\}$ sampled from $\mathcal{D}=\{e_{1},e_{2},\dots e_{N}\}$. The loss function is optimized by taking the gradient with respect function approximation weights
$\theta_{i+1}=\theta_{i}+\alpha(r(s_{t},a_{t},s_{t+1})+\gamma\mathcal{Q}(s_{t+1},$ $\operatorname*{arg\,max}_{a}\mathcal{Q}(s_{t+1},a;\theta^{\textrm{target}}_{i-1});\theta^{\textrm{target}}_{i-1})$
$\qquad\qquad\qquad-\mathcal{Q}(s_{t},a_{t};\theta_{i}))\nabla_{\theta_{i}}\mathcal{Q}(s_{t},a_{t};\theta_{i}).$
In a parallel line of algorithm families the policy itself is directly parametrized by $\pi_{\theta}$ (Sutton et al.,, 1999), and the gradient estimator used in learning is
@@
where $Q(s_{t},a_{t})$ refers to the state-action value function at time step $t$. The algorithms that focus on directly parameterizing the policy try to solve the following optimization problem (Schulman et al.,, 2015).
$\max_{\theta}\mathbb{E}_{s\sim\rho_{\theta_{\textrm{old}}};a\sim\pi_{\theta_{\textrm{old}}}(s,\cdot)}\left[\dfrac{\pi_{\theta}(s,a)}{\pi_{\theta_{\textrm{old}}}(s,a)}Q_{\theta_{\textrm{old}}}(s,a)\right]\>\>\textrm{subject to}\>\>$ $\mathbb{E}_{s\sim\rho_{\theta_{\textrm{old}}}}\mathcal{D}_{KL}(\pi_{\theta}(s,\cdot)||\pi_{\textrm{old}}(s,\cdot))\leq\delta$
@@ 3.1 Generic Reinforcement Learning Algorithm
To be able to understand and analyze the connection between different approaches to achieve generalization first we will provide a clear definition intended to capture the behavior of a generic reinforcement learning algorithm.
Definition 3.1 (Generic reinforcement learning algorithm).
@@
A reinforcement learning training algorithm $\mathcal{A}$ learns a policy $\pi$ by interacting with an MDP $\mathcal{M}$. We divide up the execution of $\mathcal{A}$ into discrete time steps as follows. At each time $t$, the algorithm has a current policy $\pi_{t}$, observes a state $s_{t}$, takes an action $a_{t}\sim\pi_{t}(s_{t},\cdot)$, and observes a transition to state $s^{\prime}_{t}\sim\mathcal{P}(\cdot\mid s_{t},a_{t})$ with corresponding reward $r_{t}=r(s_{t},a_{t},s^{\prime}_{t})$. We define the history of algorithm $\mathcal{A}$ in MDP $\mathcal{M}$ to be the sequence $H_{t}=(\pi_{0},s_{0},a_{0},s^{\prime}_{0},r_{0}),\dots(\pi_{t},s_{t},a_{t},s^{\prime}_{t},r_{t})$ of all the transitions observed by the algorithm so far. We require that the policy $\pi_{t}$ and state $s_{t}$ at time $t$ are a function only of $H_{t-1}$, i.e the transitions observed so far by $\mathcal{A}$. At time $t=T$, the algorithm stops and outputs the policy $\pi=\pi_{T}$. We use the notation $\mathbb{A}$ to denote the set of reinforcement learning training algorithms and $\Pi$ to denote the set of policies $\pi$ in an MDP $\mathcal{M}$.
@@
Intuitively, a reinforcement learning algorithm has a current policy $\pi_{t}$, performs a sequence of queries $(s_{t},a_{t})$ to the MDP, and observes the resulting state transitions and rewards. In order to be as generic as possible, the definition makes no assumptions about how the algorithm chooses the sequence of queries, other than that $a_{t}\sim\pi_{t}(s_{t},\cdot)$. Notably, if taking action $a_{t}$ in state $s_{t}$ leads to a transition to state $s^{\prime}_{t}$, there is no requirement that $s_{t+1}=s^{\prime}_{t}$. Indeed, the only assumption is that $s_{t+1}$ and $\pi_{t+1}$ may depend only on $H_{t}$, the history of transitions observed so far. This allows the definition to capture deep reinforcement learning algorithms, which may choose to query states and actions in a complex way as a function of previously observed state transitions.
3.2 Base Generalization in Deep Reinforcement Learning
We next introduce a basic metric capturing how well an algorithm generalizes given a fixed amount of interaction with a given MDP.
Definition 3.2 (Base generalization).
Given an MDP $\mathcal{M}=(S,A,P,r,\rho_{0},\gamma)$, let $\pi_{T}$ and $\hat{\pi}_{T}$ be policies output by training algorithms taking $T$ steps. The base generalization $\mathcal{G}^{\textrm{base}}$ is the difference between the expected discounted cumulative rewards obtained by policy $\pi_{T}$ and $\hat{\pi}_{T}$ in $\mathcal{M}$.
$\mathcal{G}^{\textrm{base}}(\pi_{T},\hat{\pi}_{T})=\mathbb{E}_{a_{t}\sim\pi_{T}(s_{t},\cdot)}$ $\left[\sum_{t=0}^{\infty}\gamma^{t}r(s_{t},a_{t},s_{t+1})\right]$
$\qquad\quad-\mathbb{E}_{\hat{a}_{t}\sim\hat{\pi}_{T}(\hat{s}_{t},\cdot)}\left[\sum_{t=0}^{\infty}\gamma^{t}r(\hat{s}_{t},\hat{a}_{t},\hat{s}_{t+1})\right]$
The base generalization definition captures how well an algorithm can generalize to unseen states and transitions, given only access to $T$ interactions with the MDP $\mathcal{M}$. Hence, in base generalization the role of exploration is exceedingly dominant and this will be further explained in Section 5.
3.3 Algorithmic Generalization
Based on the definition of generic reinforcement learning algorithm, we will now further define the different approaches proposed to achieve generalization. At a high level, the approaches we will discuss will be divided into two classes: I. Techniques that solely modify the training algorithm,
II. Techniques that directly modify the MDP (i.e. learning environment, training data) that forms the interactions of the training algorithm with the learning environment.
Our first definition formalizes the techniques that solely modify the training algorithm.
Definition 3.3 (Algorithmic generalization).
Let $\mathcal{A}$ be a training algorithm that takes an MDP as input and outputs a policy. Given an MDP $\mathcal{M}=(S,A,P,r,\rho_{0},\gamma)$, an algorithmic generalization method $\mathcal{G}_{\mathbb{A}}$ is given by a function $F:\mathbb{A}\to\mathbb{A}$ that runs the algorithm $F(\mathcal{A})$ in the MDP $\mathcal{M}$.
Algorithmic generalization captures modifications to the training algorithm itself that can range from the choice of optimization methods or regularizers, to update rules for the policy.
3.4 Generalization Through Rewards
@@
Let $\mathcal{A}$ be a training algorithm that takes as input an MDP and outputs a policy. Given an MDP $\mathcal{M}=(S,A,P,r,\rho_{0},\gamma)$, a rewards transforming generalization method $\mathcal{G}_{R}$ is given by a sequence of functions $F_{t}:(\Pi\times S\times A\times S\times\mathbb{R})^{t}\times\mathbb{R}\to\mathbb{R}$. The method attempts to achieve generalization by running $\mathcal{A}$ on MDP $\mathcal{M}$, but modifying the rewards at each time $t$ to be $\hat{r}_{t}(s_{t},a_{t},s^{\prime}_{t})=F_{t-1}(H_{t-1},r_{t})$, where $H_{t-1}$ is the history of algorithm $\mathcal{A}$ when running with the transformed rewards.
In particular, a method under the rewards transforming generalization category runs the original algorithm to train the policy, but modifies the observed rewards. The instances of these techniques will be mentioned and explained in Section 6.2, i.e. direct function regularization, in Section 5, i.e. the role of exploration in overfitting, and in Section 8, i.e. transfer in reinforcement learning.
3.5 Generalization Through Observations
Following the definition of reward transforming generalization we define state transforming generalization which is one of the canonical approaches for achieving generalization in deep reinforcement learning. The instances of generalization through observations will be categorized and explained in detail in Section 6.1, i.e. data augmentation, and Section LABEL:adv, i.e. the adversarial perspective for deep neural policy generalization.
@@ 3.5 Generalization Through Observations
Let $\mathcal{A}$ be a training algorithm that takes as input an MDP and outputs a policy. Given an MDP $\mathcal{M}=(S,A,P,r,\rho_{0},\gamma)$, a state transforming generalization method $\mathcal{G}_{S}$ is given by a sequence of functions $F_{t}:(\Pi\times S\times A\times S\times\mathbb{R})^{t}\times S\to S$. The method attempts to achieve generalization by running $\mathcal{A}$ on MDP $\mathcal{M}$, but modifying the state chosen at time $t$ to be $\hat{s}_{t}=F_{t-1}(H_{t-1},s_{t})$, where $H_{t-1}$ is the history of algorithm $\mathcal{A}$ when running with the transformed states.
3.6 Generalization Through Environment Dynamics
Another category of algorithms that tries to achieve generalization in deep reinforcement learning focuses on achieving this objective through environment dynamics transformation. The methods focusing on generalization through environment dynamics will be referred to and explained in Section 6.2, i.e. direct function regularization.
@@ 3.6 Generalization Through Environment Dynamics
Let $\mathcal{A}$ be a training algorithm that takes as input an MDP and outputs a policy. Given an MDP $\mathcal{M}=(S,A,P,r,\rho_{0},\gamma)$, a transition probability transforming generalization method $\mathcal{G}_{\mathcal{P}}$ is given by a sequence of functions $F_{t}:(\Pi\times S\times A\times S\times\mathbb{R})^{t}\times(S\times A\times S)\to\mathbb{R}$. The method attempts to achieve generalization by running $\mathcal{A}$ on MDP $\mathcal{M}$, but modifying the transition probabilities at time $t$ to be $\hat{P}(s_{t},a_{t},s^{\prime}_{t})=F_{t-1}(H_{t-1},s_{t},a_{t},s^{\prime}_{t})$, where $H_{t-1}$ is the history of algorithm $\mathcal{A}$ when running with the transformed transition probabilities.
3.7 Generalization Through Policy
The last type of generalization method we define is based on directly modifying the current policy used by the algorithm to select actions at each time step. We will explain the instances of the techniques that focus on generalization through policy in Section 5, i.e. the role of exploration in overfitting, and Section 7, i.e. meta reinforcement learning and meta gradients.